1/18h^2-1/2h+1=0
Solve each equation- I am not sure how to get started on this problem. HELP Please

Answer by ankor@dixie-net.com(22740)

1/18h^2 - 1/2h + 1 = 0
:
First get rid of those denominators, multiply equation by 18 and you have:
1h^2 - 9h + 18 = 0
:
Factors easily to:
(h - 3)(h - 6) = 0
h = +3
and
h = +6
:
Check our solutions: substitute 3 for h in the original equation:
(1/18)3^2 - (1/2)3 + 1 =
(1/8)9 - 3/2 + 1`=
9/18 - 3/2 + 1 =
1/2 - 3/2 + 1 = 0
:
I'll let you check the h = 6 solution